A solutian is made by mixing 32.0 ml of ethanol cdHo, and 68 o mL of water. Assuming ideal behavior, what is the vapor pressure of the solution at 20 C? Values at 20 C Density Vapor Pressure, P Liquid Torr 2 Check answer Nest Eet
Answer #1

we can do this via molar fractions

so..

mol of ethanol = mass/MW

and mass = D*V so

mol = D*V/MW = (0.789*32)/(46) = 0.5488

for

water:

mol = D*V/MW = (0.998*68)/(18) = 3.770

now...

total mol = 0.5488+3.770 = 4.3188

Pvapor mix = x1*P1° + x2*P°2 = (3.770)/4.3188*17.5 + 0.5488/4.3188*43.9

Pvapor mix = 20.85 torr